A student adds 207.6 grams of C7H6O3 to an excess of C4H6O3, this produces C9H8O4 and C2H4O2. Calculate the percent yield if 231 grams of aspirin (C9H8O4) is produced


1
Expert's answer
2021-02-05T04:05:21-0500

C7H6O3 + C4H6O3 = C9H8O4 + C2H4O2

mtheoretical (C9H8O4) = m(C7H6O3 ) × Mr(C9H8O4 ) / Mr(C7H6O3) = 207.6 g × 180 g/mol / 138 g/mol = 270.8 g

%w = (mactual / mtheoretical) × 100 % = (231 g / 270.8 g) × 100 % = 85.3 %


Answer: 85.3 %